# Variational Prior Replacement in Bayesian Inference and Inversion Xuebin Zhao1 and Andrew Curtis1

### v1

@@ 3.2 Verifying variational prior replacement
In the first test, we verify that VPR produces correct results by comparing them to those obtained using a conventional approach where an independent Bayesian inversion is performed for each prior distribution (referred to herein as prior specific inversion). For this test, we consider the Uniform prior distribution and the smoothed prior . For the prior specific case, PSVI is used to solve these two FWI problems with the same respective priors. We update variational parameters (mean vector and lower triangular matrix mentioned in Section 2.5) for 5,000 iterations. During each iteration, 2 random samples are employed to approximate the ELBO (equation 8) using Monte Carlo integration; such a low number of samples has been shown to be reasonable in a stochastic sense in previous studies, because of the large number of iterations. For VPR, the Uniform distribution is treated as the old prior distribution, which is then removed from the old posterior distribution (the posterior pdf calculated using by prior specific inversion) and replaced by the smoothed (new) prior distribution . This is achieved by solving the variational problem described in equation 13, through minimisation of the KL divergence expressed in equation 12. Similarly to the prior specific case, PSVI is used to solve this problem, where variational parameters are updated for 5,000 iterations with 10 samples per iteration used to estimate the expectation term in equation 12.
@@
Figures 5a and 5b display prior specific inversion (PSI) and VPR results obtained using the smoothed prior distribution. In each column, a random posterior sample, the mean velocity map, standard deviation and the relative error of the posterior distribution are presented from top to bottom row. The relative error is calculated as the difference between the mean and true velocity models (Figure 1a) divided by the standard deviation at each point, reflecting the relative deviation between the true and inverted mean models. The most important feature is that the first-order posterior statistics displayed in Figures 5a and 5b are almost identical; Figure 5b was produced from the results obtained using the Uniform prior (displayed in Figure 8a), which were significantly different. This proves that VPR can accurately replace the old prior information and inject (update) the new prior information into the inversion results, without solving a Bayesian inverse problem again from scratch.
@@
We do observe some discrepancies between these two sets of results. For example, a vertically oriented low velocity structure is present in Figure 5a inside dashed black and red boxes, which is not present in Figure 5b. Although this feature is not observed in the true velocity model in Figure 1a, there is no strong evidence to discriminate which result is better. Nonetheless, it is reasonable to have these discrepancies, which are caused by the fact that in VPR we introduce a variational distribution to approximate the new posterior distribution , as expressed in equation 13, rather than calculating the actual (Walker & Curtis, 2014b).
@@
Figure 6 compares the posterior marginal pdfs of the above two results along two vertical velocity profiles at horizontal locations of 1km (top row) and 2.6km (bottom row). Their locations are displayed by dashed black lines in Figure 1a. Red and black lines represent the true and mean velocity values, respectively. Despite some small discrepancies presented here and there in Figures 6a and 6b, they provide similar posterior marginal pdfs. Interestingly, as displayed in the top row, the two methods (PSI and VPR) find very similar yet incorrect posterior solutions given the same data and prior information (especially below 1.3km depth where true velocity values are excluded by the high probability region of the posterior pdfs). This is because that the PSVI algorithm might find a poor solution at this region caused by cycle skipping, which occurs typically in FWI problems. We also compare correlation information of the two posterior pdfs in Figure 7, which displays the posterior correlation matrices for velocity values in a 2D window with a size of 10 10 cells inside the black box in Figure 1a. The top row shows the full correlation matrices (with a size of 100 100), and the bottom row shows the first 30 30 elements. Highly consistent posterior correlation values are obtained, which again proves the effectiveness of VPR.
@@
In conclusion, since variational prior replacement and prior specific inversion provide almost identical posterior random samples, first-order statistics (posterior mean, standard deviation and marginal pdfs) and second-order statistics (correlation matrices), we assert that the proposed method is effective and accurate for varying prior information in Bayesian inference without performing repeated independent inversions.
@@ 3.3 FWI using different priors
In this section we analyse the effect of the three different priors defined previously and compare the corresponding inversion results. We employ PSVI to perform a single variational Bayesian FWI using the Uniform distribution . is then replaced by both the smoothed prior and the geological prior using VPR. Note that in VPR the old prior distribution should be broader than the new prior to avoid numerical instability issues such as occurs when attempting to divide by zero (Walker & Curtis, 2014b). However, the Gaussian geological prior distribution is defined in the space of real numbers, which spans a broader parameter space than the Uniform distribution. Therefore, we truncate within the lower and upper bounds of the Uniform prior distribution, and renormalise by another normalisation constant (which does not need to be evaluated, similar to those in equations 20 and 22).
Figures 8a – 8c display the obtained inversion results. Each figure includes a random posterior sample, the mean velocity, standard deviation and the relative error maps of the posterior pdf from top to bottom row. Note that Figures 5b and 8b represent the same results obtained using VPR. In a previous study (Zhao & Curtis, 2024b), we compared the inversion results obtained without using prior replacement displayed in Figure 8a with two entirely independent variational methods using exactly the same Uniform prior distribution and observed data and obtained highly consistent results, proving that we obtain approximately correct posterior uncertainty statistics for this specific prior. In this study we focus on the inversion results obtained using different priors.
The posterior random sample in Figure 8a shows significant velocity ‘speckle’ - strong, short wavelength contrasts - since no correlation is introduced by the Uniform prior distribution. The relatively non-informative prior information results in significantly higher uncertainties at greater depths, increasing up to around 800 m/s. The two posterior samples in Figures 8b and 8c are smoother since extra (smooth) prior information is injected into the two inversion results which precludes sharp velocity variations between neighbouring cells. The smoothed prior pdf imposes spatial smoothness explicitly, and the geological prior pdf injects similar information implicitly, as illustrated by the positive correlations displayed in Figure 3. Therefore, the true velocity structures are better resolved since they are indeed laterally fairly smooth.
@@
Standard deviation values displayed in Figures 8b and 8c are smaller than those in Figure 8a (note that different colorbars are used in these figures). This makes sense because by imposing prior information that velocity structures should be relatively smooth, we have removed the possibility of including large velocity contrasts between laterally proximal cells. The posterior uncertainties in Figure 8c are higher than those in Figure 8b, possibly due to the different magnitude of smoothness applied to these two results (one controlled by the predefined parameter in equation 19 and the other by correlations calculated from the training images such as Figure 2). In addition to the magnitude of the standard deviation values, strong prior information also suppresses overall changes in the uncertainty structures: obvious spatial (vertical) variations are observed in the standard deviation map in Figure 8a. For example, uncertainties increase at depth since data sensitivity decreases at depth and lower uncertainties are observed around some high velocity layers. However, those features are less significant in Figure 8c and almost invisible in Figure 8b. This is because the geological prior information imposes weaker vertical smoothness, as illustrated in Figures 3 and 4c, whereas the smoothed prior distribution imposes the same magnitude of smoothness in both vertical and horizontal directions. As a result, we observe larger relative errors in Figures 8b and 8c, especially at layer boundaries where higher uncertainties should be expected (Galetti et al., 2015).
@@
Figure 9 displays the posterior marginal pdfs of the three inversion results at the same two locations as in Figure 6. Similarly to the standard deviation maps, the marginal pdfs in Figures 9b and 9c are narrower than those in Figure 9a due to the additional prior information injected. Figure 9c presents larger vertical variations than Figure 9b. As marked by the three white arrows in the first row, the (old) posterior pdf using the Uniform prior distribution fails to find the true solution, thus neither do the two VPR results. This is reasonable considering the variational prior replacement methodology: the method can only replace prior information imposed previously into the inversion results, but it cannot correct the old estimate of the posterior distribution in cases where the estimate is poor. Nevertheless, in places where the old posterior distribution is correct, VPR injects new prior information correctly as displayed in the bottom row.
@@
To analyse posterior correlation information in both horizontal and vertical directions, we calculate the correlation matrices for velocity values selected from 10 horizontally contiguous and 10 vertically contiguous grid cells (marked by top and left boundaries of the black box in Figure 1a). The corresponding results are displayed in the top and bottom rows of Figure 10. Similarly to Figures 8 and 9, each column (Figures 10a – 10c) represents the calculated correlation matrix from one specific prior pdf. Since no correlation information is introduced by the Uniform prior distribution, the posterior correlations displayed in Figure 10a are purely determined by observed data. Negative correlations are observed between neighbouring cells and positive correlations are vaguely presented between every second neighbouring cells. Since FWI is a highly underdetermined inverse problem (the effective number of independent data points is significantly smaller than the number of unknown model parameters), velocity values oscillate between adjacent cells to achieve a better data fit, especially for grid cells within one wavelength. Similar posterior correlation patterns using a Uniform prior distribution were observed in previous studies (Gebraad et al., 2020; Zhang et al., 2023; Zhao & Curtis, 2024b). This would also happen in linearised (deterministic) FWI if no regularisation term (prior information) was added.
@@ 4 Discussion
We demonstrated that variational prior replacement (VPR) can change prior information efficiently post Bayesian inference. The updated posterior distribution is found by solving a variational problem, in which a variational distribution is introduced and optimised iteratively to approximate , as expressed in equation 13. Therefore, we do not expect VPR and prior specific inversion (PSI) to provide exactly the same results, especially if PSI itself is performed using variational inference as in this study which also results in an approximation. This is the reason why we observe some small discrepancies between the results obtained using VPR and PSI displayed in Figures 5 – 7.
@@
In addition, (semi-)analytic calculation of equation 5 requires the evaluation of the normalisation constant by integrating , and over the entire parameter space analytically, which might be intractable for high dimensional inference problems, and indeed only under certain circumstances can this be done (Walker & Curtis, 2014b). In the proposed VPR framework, we introduce a second variational distribution , which is found by minimising the KL-divergence in equation 12; the specific value of then need not be estimated explicitly, so VPR can be implemented in a more straightforward manner.
One essential requirement for prior replacement developed here and in Walker & Curtis (2014b) is that the probability value for the old posterior pdf must be able to be evaluated cheaply (otherwise, there is no reason to use VPR instead of prior specific inversion). In this study, we use physically structured variational inference for this purpose (Zhao & Curtis, 2024b), which constructs a transformed Gaussian distribution with a specific correlation structure to approximate the old posterior distribution so that the probability value can be calculated efficiently. Other well-established parametric variational inference methods, such as normalising flows (Rezende & Mohamed, 2015; Dinh et al., 2015; Kobyzev et al., 2019; Papamakarios et al., 2019; Zhao et al., 2022a; Siahkoohi et al., 2020; Levy et al., 2022), automatic differentiation variational inference (Kucukelbir et al., 2017; Zhang & Curtis, 2020; Bates et al., 2022; Sun et al., 2023) and boosting variational inference (Guo et al., 2016; Miller et al., 2017; Locatello et al., 2018; Zhao & Curtis, 2024a), can also be used to construct . The choice of method should be based on the specific problem at hand, since the No Free Lunch theorem states that no method is better than any other when averaged across all problems (Wolpert & Macready, 1997).

### v2

@@ 3.2 Verifying variational prior replacement
In the first test, we verify that VPR produces correct results by comparing them to those obtained using a conventional approach where an independent Bayesian inversion is performed for each prior distribution (referred to herein as prior specific inversion). For this test, we consider the uniform prior distribution and the smoothed prior . For the prior specific case, PSVI is used to solve these two FWI problems with their respective priors. We update variational parameters (mean vector and lower triangular matrix mentioned in Section 2.5) for 5,000 iterations. During each iteration, 2 random samples are employed to approximate the ELBO (equation 8) using Monte Carlo integration; such a low number of samples has been shown to be reasonable in a stochastic sense in previous studies, because of the large number of iterations (Kucukelbir et al., 2017). For VPR, the uniform distribution is treated as the old prior distribution, which is then removed from the old posterior distribution (the posterior pdf calculated using by prior specific inversion) and replaced by the smoothed (new) prior distribution . This is achieved by solving the variational problem described in equation 13, through minimisation of the KL divergence expressed in equation 12. Similarly to the prior specific case, PSVI is used to solve this problem, where variational parameters are updated for 5,000 iterations with 10 samples per iteration used to estimate the expectation term in equation 12. Note that it might be impossible to replace a smoothed (old) prior by a Uniform (new) prior distribution using VPR since in this case the support of the old prior may only be a subset of that of the new prior distribution (or it might be effectively so due to sampling and numerical approximations). This might make numerically unstable because for some parameters the value could be small, poorly determined or even zero.
@@
Figures 5a and 5b display prior specific inversion (PSI) and VPR results obtained using the smoothed prior distribution. In each column, a random posterior sample, the mean velocity map, standard deviation and the relative error of the posterior distribution are presented from top to bottom row. The relative error defined to be the difference between the mean and true velocity models (Figure 1a) divided by the standard deviation at each point, reflecting the relative deviation between the true and inverted mean models. The most important feature is that the first-order posterior statistics displayed in Figures 5a and 5b are almost identical; Figure 5b was produced from the results obtained using the uniform prior (displayed in Figure 9a). This supports the statement that VPR is able to replace the old prior information and inject (update) the new prior information into the inversion results, without solving a Bayesian inverse problem again from scratch.
@@
We do observe some discrepancies between these two sets of results. For example, a vertically oriented low velocity structure is present in Figure 5a inside dashed black and red boxes, which is not present in Figure 5b. Although this feature is not observed in the true velocity model in Figure 1a, there is no strong evidence to discriminate which result is better.
For other local regions such as those below and to the right of the boxes in Figure 5, the mean velocity model from VPR (Figure 5b) appears to be closer to the old posterior distribution using the uniform prior pdf (displayed in Figure 9a) than it does to the PSI result using the smoothed prior (Figure 5a). This might be because this implementation of PSI (which is an optimisation problem) converged to a local minimum, or that it has not fully converged (full convergence might require a larger number of forward evaluations which is very expensive). Alternatively, it is also possible that the VPR procedure (also an optimisation process) has not fully converged and thus might indicate an incomplete prior replacement in this test. On the other hand, for some other statistics such as the characteristic of spatial variations in each posterior sample, standard deviations, or posterior marginal pdfs displayed in Figure 6, VPR results are clearly closer to the PSI results using the smoothed prior than to those using the uniform prior, supporting the statement that VPR produces reasonable statistical accuracy. We also note that it is reasonable that there remain some discrepancies, caused by the fact that in VPR we introduce a variational distribution to approximate the new posterior distribution as expressed in equation 13, rather than calculating the actual as in Walker & Curtis (2014b).
@@
Figure 6 compares the posterior marginal pdfs of the above two results along two vertical velocity profiles at horizontal locations of 1 km (top row) and 2.6 km (bottom row). Their locations are displayed by dashed black lines in Figure 1a. Red and black lines represent the true and mean velocity values, respectively. Despite some small discrepancies here and there in Figures 6a and 6b, they provide similar posterior marginal pdfs. Interestingly, as displayed in the top row, the two methods (PSI and VPR) find very similar yet incorrect posterior solutions given the same data and prior information (especially below 1.3 km depth where true velocity values are excluded by the high probability region of the posterior pdfs). This is because that the PSVI algorithm may have converged around an incorrect solution in this region caused by cycle skipping, which often occurs in FWI problems. We also compare correlation information of the two posterior pdfs in Figure 7, which displays the posterior correlation matrices for velocity values in a 2D window with a size of 10 10 cells inside the black box in Figure 1a. The top row shows the full correlation matrices (with a size of 100 100), and the bottom row shows the first 30 30 elements. Highly consistent posterior correlation values are obtained, which again proves the effectiveness of VPR.
To further test the performance of VPR, in Figures 8a and 8b we compare one observed shot gather with data simulated by one randomly chosen posterior sample obtained from PSI and VPR results, respectively. In both figures, the simulated data are highly consistent with the observed (noisy) data, which demonstrates that VPR can provide models that produce accurate waveform data that fit the observed data to within data uncertainties.
@@
In Appendix A, we present a second example to test the accuracy of VPR using a Gaussian prior distribution for , which again shows that VPR and PSI provide highly consistent inversion results and posterior statistics. In conclusion, since variational prior replacement and prior specific inversion provide almost identical posterior random samples, first-order statistics (posterior mean, standard deviation and marginal pdfs) and second-order statistics (correlation matrices), we assert that the proposed method is effective and accurate for varying prior information in Bayesian inference without performing repeated independent inversions.
@@ 3.3 FWI using different priors
In this section we analyse the effect of the three different priors defined previously and compare the corresponding inversion results. We employ PSVI to perform a single variational Bayesian FWI using the uniform distribution . is then replaced by both the smoothed prior and the geological prior using VPR. Note that in VPR the old prior distribution should be broader than the new prior to avoid numerical instability issues such as occur when attempting to divide by zero (Walker & Curtis, 2014b). However, the Gaussian geological prior distribution is defined in the space of real numbers, which spans a broader parameter space than the uniform distribution. Therefore, we truncate within the lower and upper bounds of the uniform prior distribution, and renormalise by another normalisation constant (which does not need to be evaluated, similar to those in equations 20 and 22).
Figures 9a – 9c display the obtained inversion results. Each figure includes a random posterior sample, the mean velocity, standard deviation and the relative error maps of the posterior pdf from top to bottom row. Note that Figures 5b and 9b represent the same results obtained using VPR. In a previous study (Zhao & Curtis, 2024b), we compared the inversion results obtained without using prior replacement displayed in Figure 9a with two entirely independent variational methods using exactly the same uniform prior distribution and observed data and obtained highly consistent results, proving that we obtain approximately correct posterior uncertainty statistics for this specific prior. In this study we focus on the inversion results obtained using different priors.
The posterior random sample in Figure 9a shows significant velocity ‘speckle’ - strong, short wavelength contrasts - since no correlation is introduced by the uniform prior distribution. The relatively non-informative prior information results in significantly higher uncertainties at greater depths, increasing up to around 800 m/s. The two posterior samples in Figures 9b and 9c are smoother since extra (smooth) prior information is injected into the two inversion results which precludes sharp velocity variations between neighbouring cells. The smoothed prior pdf imposes spatial smoothness explicitly, and the geological prior pdf injects similar information implicitly, as illustrated by the positive correlations displayed in Figure 3. Therefore, the true velocity structures are better resolved since they are indeed laterally fairly smooth.
@@
Standard deviation values displayed in Figures 9b and 9c are smaller than those in Figure 9a (note that different colorbars are used in the former figures). This makes sense because by imposing prior information that velocity structures should be relatively smooth, we have removed the possibility of including large velocity contrasts between laterally proximal cells. In fact, the introduction of prior information, if it correctly reflects the true state of nature, should lead to models that better reflect the true state of nature overall (other than in pathalogical cases). In our case, the geological prior information introduced is derived from pictures that represent real geology and is therefore reasonably reflective of the true model, so in this case at least, the introduction of prior information should improve the result. Nevertheless, in some cases the uncertainty reduction displayed in Figures 9b and 9c might not be a good outcome. Normally there is less information at greater depths from waveform data recorded at the surface. The decreased uncertainties at depth occur because information from data and prior knowledge are combined. This only leads to more accurate models if the prior information is also accurate. In the case of our smoothed prior, smoothing is applied equally in vertical and horizontal directions which does not correctly reflect the structure of the true model. In that case we therefore would not expect that results are more accurate after applying this prior information (but they may be, for example if the relative errors in vertical smoothing are more than compensated by increased accuracy in horizontal smoothing). In addition, the posterior uncertainties in Figure 9c are higher than those in Figure 9b, possibly due to the different magnitude of smoothness applied to these two results (one controlled by the predefined parameter in equation 19 and the other by correlations calculated from the training images such as Figure 2).
In addition to the magnitude of the standard deviation values, strong prior information also suppresses overall changes in the uncertainty structures: obvious spatial (vertical) variations are observed in the standard deviation map in Figure 9a. For example, uncertainties increase at depth since data sensitivity decreases at depth and lower uncertainties are observed around some high velocity layers. However, those features are less significant in Figure 9c and almost invisible in Figure 9b. This is because the geological prior information imposes weaker vertical smoothness, as illustrated in Figures 3 and 4c, whereas the smoothed prior distribution imposes the same magnitude of smoothness in both vertical and horizontal directions. As a result, we observe larger relative errors in Figures 9b and 9c, especially at layer boundaries where higher uncertainties should be expected (Galetti et al., 2015).
@@
Figure 10 displays the posterior marginal pdfs of the three inversion results at the same two locations as in Figure 6. Similarly to the standard deviation maps, the marginal pdfs in Figures 10b and 10c are narrower than those in Figure 10a due to the additional prior information injected. Figure 10c presents larger vertical variations than Figure 10b. As marked by the three white arrows in the first row, the (old) posterior pdf using the uniform prior distribution fails to find the true solution, thus neither do the two VPR results. This is reasonable considering the variational prior replacement methodology: the method can only replace prior information imposed previously into the inversion results, but it cannot correct (improve) the old estimate of the posterior distribution in cases where the old estimate is poor. Nevertheless, in places where the old posterior distribution includes the true model solution, VPR injects new prior information properly as displayed in the bottom row.
@@
To analyse posterior correlation information in both horizontal and vertical directions, we calculate the correlation matrices for velocity values selected from 10 horizontally contiguous and 10 vertically contiguous grid cells (marked by top and left boundaries of the black box in Figure 1a). The corresponding results are displayed in the top and bottom rows of Figure 11. Similarly to Figures 9 and 10, each column (Figures 11a – 11c) represents the calculated correlation matrix from one specific prior pdf. Since no correlation information is introduced by the uniform prior distribution, the posterior correlations displayed in Figure 11a are purely determined by observed data. Negative correlations are observed between neighbouring cells and positive correlations are vaguely presented between every second neighbouring cells. Since FWI is a highly underdetermined inverse problem (the effective number of independent data points is significantly smaller than the number of unknown model parameters), velocity values oscillate between adjacent cells to achieve a better data fit, especially for grid cells within one wavelength. Similar posterior correlation patterns using a uniform prior distribution were observed in previous studies (Gebraad et al., 2020; Zhang et al., 2023; Zhao & Curtis, 2024b). This would also happen in linearised (deterministic) FWI if no regularisation term (prior information) was added.
@@ 4 Discussion
We demonstrated that variational prior replacement (VPR) can change prior information efficiently post Bayesian inference. The updated posterior distribution is found by solving a variational problem, in which a variational distribution is introduced and optimised iteratively to approximate , as expressed in equation 13. Therefore, we do not expect VPR and prior specific inversion (PSI) to provide exactly the same results, especially if PSI itself is performed using variational inference as in this study which also results in an approximation. This is part of the reason why we observe some small discrepancies between the results obtained using VPR and PSI displayed in Figures 5 – 7.
@@
In addition, (semi-)analytic calculation of equation 5 requires the evaluation of the normalisation constant by integrating , and over the entire parameter space analytically, which might be intractable for high dimensional inference problems, and indeed only under certain circumstances can this be done (Walker & Curtis, 2014b). In the proposed VPR framework, we introduce a second variational distribution , which is found by minimising the KL-divergence in equation 12; the specific value of then need not be estimated explicitly, so VPR can be implemented in a more straightforward manner. On the other hand, this implies that VPR is itself an approximate method which uses to approximate .
One essential requirement for prior replacement developed here and in Walker & Curtis (2014b) is that the probability value for the old posterior pdf must be able to be evaluated cheaply (otherwise, there is no reason to use VPR instead of prior specific inversion). In this study, we use physically structured variational inference for this purpose (Zhao & Curtis, 2024b), which constructs a transformed Gaussian distribution with a specific correlation structure to approximate the old posterior distribution so that the probability value can be calculated efficiently. Other well-established parametric variational inference methods, such as normalising flows (Rezende & Mohamed, 2015; Dinh et al., 2015; Kobyzev et al., 2020; Papamakarios et al., 2021; Zhao et al., 2022a; Siahkoohi et al., 2020; Levy et al., 2022), automatic differentiation variational inference (Kucukelbir et al., 2017; Zhang & Curtis, 2020a; Bates et al., 2022; Sun et al., 2023) and boosting variational inference (Guo et al., 2016; Miller et al., 2017; Locatello et al., 2018; Zhao & Curtis, 2024a), can also be used to construct . The choice of method should be based on the specific problem at hand, since the No Free Lunch theorem states that no method is better than any other when averaged across all problems (Wolpert & Macready, 1997).
@@
Based on a Markov chain Monte Carlo (McMC) framework, Mosegaard & Tarantola (1995) introduced an approach to draw samples from one probability distribution when one only has samples from another distribution. They achieved this by resampling using a simple Metropolis accept-reject criterion based on the ratio of two probability values , potentially avoiding further likelihood function evaluations. In this sense, that method can be interpreted as a Monte Carlo version of prior replacement that provides a sampling-based solution, compared to the VPR approach which is based on variational inference. Both approaches of Mosegaard & Tarantola (1995) and in this paper have a similar pre-requisite: that the support of the old prior distribution should include that of the new prior so that numerical instability issues are avoided when performing the division .
@@
The main purpose of VPR is to update (replace) prior information efficiently in Bayesian inference. As demonstrated herein, new inversion results can be obtained with no further forward simulation. On the other hand, this also indicates that we cannot obtain new information about data misfit values purely from VPR results - indeed, that is the point of the method. In future, VPR might be used to compare different prior hypotheses (e.g., including different magnitudes of smoothness for a smoothed prior distribution): say we have obtained approximate posterior pdfs for a set of different prior assumptions by applying VPR. For the different inversion results, we could perform a small number of additional forward simulations using posterior samples drawn from VPR results, based on which the different prior hypotheses can be compared, and one or two close-to-optimal options could be selected. Although this procedure does require additional forward simulations, it still provides a far more efficient approach to test different prior options compared to carrying out a sequence of independent inversions as is typically done using linearised inversion. In addition, for cases when VPR becomes less accurate (e.g., when the dimensionality and complexity of the inverse problem increase such as in 3D FWI problems), one might use a relatively lower cost forward function with different data types to refine (fine tune) the outcomes obtained from VPR, by invoking Bayes’ rule again.